# Autonomous multi-ion optical clock with on-chip integrated photonic light delivery

### v2

@@ I Introduction
Optical atomic clocks, as time and frequency standards, offer unparalleled precision for applications in fundamental physics [RevModPhys87637, PhysRevLett116063001, Safronova:2018, brewer:2019a, hb3cdk28, PhysRevLett134093201], navigation [Burt:2021], and telecommunications [Lindvall:25, sheridan2025telecommunications]. Neutral atom optical lattice clocks are widely used platforms that provide high precision and accuracy [Takamoto:2005, Boyd:2005, Bothwell:2019, McGrew:2018] but contain complex atom-atom interactions and require large high-power laser systems and atomic ovens. On the other hand, trapped-ion clocks [Chou:2010, brewer:2019a], provide excellent isolation of individual ions with minimal disturbance of internal states. However, they suffer from high quantum projection noise as a result of the comparatively small number of ions. Although both platforms serve as excellent clocks, they both rely on bulky optical components, which limit their operation to laboratory environments.
The path to deploying these instruments in real-world applications lies in miniaturization and ruggedization [Newman:2018, Burt:2021], a challenge for which integrated photonics provides a compelling solution. By replicating complex optical functionalities on scalable chip-based platforms [chiaverini:2005, stick:2005], this approach not only works to reduce the system’s footprint but also enables advanced multi-ion clock architectures that are less practical with conventional setups [rosenband:2013, blain:2021]. Additionally, the integration of photonic components directly into ion trap chips is a key strategy for scaling these quantum systems. Replacing complex free-space optics with on-chip waveguides and splitters [sundaram2022highly, ricci_standingwave_2023, Kwon:2023], grating couplers [Mehta2016, niffenegger:2020, mehta:2020], and photonic control [hogle:2023] provides robust and scalable delivery of the multiple laser wavelengths required for narrow transition control. These technologies are also applicable to trapped-ion quantum computing, notably the Quantum Charge-Coupled Device (QCCD) architecture [noel:2019, pino:2021], which requires multiplexed operations and coherent control across many distinct trapping zones [Mordini:2025]. By providing a pathway to individually address dense ion arrays, integrated photonics are critical for moving beyond small-scale experiments, towards large-scale quantum platforms.
Recently demonstrated building blocks for this architecture include simultaneous multi-site optical addressing [Kwon:2023] and coherent transport of ions between zones [Mordini:2025], establishing the viability of complex on-chip control primitives. Although these advances are foundational, they have so far focused on individual capabilities rather than their combination into a unified, application-focused system. Complementing these efforts in ion trap integration, there is parallel progress in developing high-stability, chip-scale lasers for integrated, high-performance light sources [Loh:2025] and integrated detectors for state readout [setzer:2021, reens:2022].
@@ II.1 Experimental setup
In previous demonstrations of integrated photonic ion traps, we utilized a free-space grating input coupling scheme [ivory:2021, Kwon:2023]. This presented challenges for both coupling efficiency and stability. Our current room-temperature system makes progress in both challenges by directly coupling an array of fibers to the trap. Single-mode, polarization-maintaining fibers in Pyrex v-groove arrays (VGAs) [Uddin:2009] were aligned with a matching array of tapered waveguides at the edge of the trap chip. The input for each beam is split on-trap via multi-mode interferometer (MMI) splitters into five waveguides, which then go to separate clock sites. Each site is addressable with four waveguide output grating windows that deliver light at 369 nm, 435 nm, 935 nm, and 760 nm. Fig.1(a) shows an illustration of the site layout on the chip. The optical loss in each waveguide path is primarily due to imperfect input coupling, material absorption, scattering during propagation, and diffraction inefficiency. The total loss for the clock beam (435 nm) exhibits a total loss of -27 dB. Although it is difficult to identify the contribution of each loss mechanism, measurements from the loop-back optimization suggest that most of the loss is due to the fiber-to-trap coupling.
For this work, we used the 435 nm E2 (clock) transition of . This additionally requires 369 nm for Doppler cooling and detection and 935 nm for repump and state preparation. A 760 nm beam can be used to pump the state to to rejoin the Doppler cooling cycle [PhysRevA104L010802], however, in our systems it does not significantly improve the overall lifetime of the ions [Kwon:2023].
@@ II.2 Clock protocol for multi-ion ensemble operation
Fig.2(a) shows the clock pulse sequence used for interrogation. Clock ions are cooled and optically pumped. Then, the ensemble of ions is probed on the E2 transition at half its spectroscopic width on either side of the inferred center for two separate second-order integrators. Probing at the half-width ensures that fluorescence counts are on average present after sufficiently many interrogations if an ion is present. Additionally, this choice gives a large derivative in the count imbalance with respect to frequency error, the basis of the error signal. The count imbalance feeds into its respective integrator to maintain reference to the ion. The integrators track each other closely as the ions are lost and reloaded at clock sites. The integrator feedback of the clock algorithm onto the local oscillator frequency is as follows. The ion is probed on the right and left sides of the transition one half-width from the inferred resonance. If the ion was shelved in the clock state and therefore below the threshold for detection, the value of the interrogation of that side is . If the ion was not shelved in the clock state, visible to the detection cycle, and therefore above the threshold for detection, the value of the interrogation of that side is . The second-order integrator () is increased by the value of the first-order integrator (). The imbalance () is then added to the first-order of the integrator. Together, these steps are as follows:
@@
In addition to managing the integrator feedback, the clock algorithm uses the same measurement to track which ions are still present and whether an ion needs to be reloaded. To determine whether the ion is present, the sum of side interrogations is tracked. Both ions that are outside of the cooling cycle and physically lost from the trap are perceived as lost when the rolling average falls below a threshold. When an ion in a clock-zone ion is lost, this is reported to the shuttling software. This report occurs every 20 clock cycles. Each clock cycle involves four interrogations: one on each side of the resonance for each integrator. One clock cycle is the distance between the dashed green lines in Fig.2(a). The reporting rate is limited by the rate at which the receiving software can communicate consistently. The remaining ions are shuttled into vacancies, automatically filling the farthest sites from the loading site first. The new ion at the loading site is stored as a buffer to quickly replenish empty clock sites. An illustration of this can be seen in Fig.2(b). The lifetime of ions at an individual clock site was correlated with the distance from the loading site. This motivated filling the farthest sites first. One hypothesis for the cause of this effect is that the localized pressure associated with neutral flux and heated surfaces increases collisions, reducing neighboring clock site lifetimes. The ion lifetimes were approximately one minute during clock operation and 15 minutes when individual ions were only being trapped and cooled. Despite this short lifetime, we were able to keep the clock sites fully occupied with four ions more than 45% of the time. Furthermore, we operated the clock system continuously for more than two hours on multiple occasions, predominantly limited by algorithmic interlocks to abort the clock in cases where the algorithm was unable to determine if a laser had become unlocked (no ions in any sites for an extended period of time). This attests to the effectiveness of the automated ion reloading, shuttling, and clock-cycle reintegration capabilities of the system.
@@ II.3 Frequency stability of integrated photonics multi-ion clock
Fig.3 shows the performance of the integrated photonics 4-ion ensemble clock. The expected quantum projection noise (QPN) of a single integrator is calculated as follows [PhysRevX14011006]:
@@ II.4 Characterization of multi-site uniformity for future scalability
The direction and focal size of each waveguide output beam is lithographically defined. Before running the clock algorithm, the ions were verified to have similar counts, Rabi frequencies, and coherence times. The electrodes must be finely tuned to match these, as small local differences are hard to predict. We also observed a dependence of the compensation voltages on the 369 nm beam power. Because the timescale of the effect was hours, the small periods of the 369 nm beam being off for the clock transition excitation were inconsequential. Once established, the compensations would hold for several weeks. While we expect to eliminate this dependence in future traps by using conductive coatings on the output grating windows, the power drift was not an issue unless the laser power was changed by roughly a factor of 2 or more. Fig.4(a) shows a representative set of Rabi flops from an ion at each site. Measurements are modeled similarly to [semenin2022determination]. The ions are trapped in separate DC potentials 180 m apart, so the chain configuration terms can be dropped, resulting in the following equation for the probability of each ion being shelved in the clock state:
@@ III Discussion
This trap suffers from high heating rates, as discussed in more detail in the Methods. Although this was not likely the limiting factor in clock performance, future systems with improved detection efficiency and ion lifetime would benefit from cooler atoms and the associated improved contrast and coherence time. One culprit for high heating rates could be noise due to the buildup of charge on the dielectric surfaces of the grating windows [PhysRevLett126230505]. This could be alleviated by the inclusion of a conductive but transparent coating over the grating windows, such as indium-tin-oxide. In addition, side-band cooling was not implemented. Side-band cooling was time prohibitive, in part because of the low optical coupling from fiber-to-trap. Better optical coupling would allow for side-band cooling and increase both contrast and coherence time.
@@ Acknowledgments.
Author Contributions.
T.D.M., J.K., D.S. and H.J.M. conceived the experiment. T.D.M. and J.K. took the measurement and analyzed data. T.D.M. developed multi-ion shuttling and clock code. M.A.D. oversaw the trap fabrication. D.R.L. advised formulation of the clock algorithm. T.D.M., J.K. and H.J.M. wrote the original manuscript. All the authors reviewed and edited the manuscript. H.J.M. supervised the project.
Competing Interests.
Author D.S. is currently employed by IonQ, Inc. The remaining authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
IV METHODS
Experimental system
@@ IV METHODS
The experiment was conducted using a trap designed and fabricated at Sandia National Laboratories. Although this trap shares features with previous designs [ivory:2021, Kwon:2023], it uniquely incorporates a multi-mode interferometer (MMI) waveguide splitter for simultaneous multi-ion clock interrogation. is a good species for deployable and robust systems due to its suppressed first-order Zeeman shift. The large wavelength range (369 nm to 935 nm) adds to the design complexity, as shorter wavelengths require unique waveguide designs and are more sensitive to fabrication tolerances and feature roughness. All waveguides were made with silicon nitride.
Fiber coupling was accomplished using a UHV vacuum-compatible glue attachment of a v-groove array (VGA). Single-mode polarization-maintaining fibers of the appropriate core diameters (to ensure single-mode propagation at the appropriate wavelength) were placed in Pyrex v-groove arrays prior to attaching [31]. These were then aligned to a matching array of tapered waveguides on the edge of the trap chip. These tapered waveguide array edge-couplers are located transverse to the trap axis to minimize the scattered light from the coupling region into the readout optics. The attachment was optimized by maximizing the light that circulates through an additional on-chip loop-back waveguide. For a fiber feedthrough, a 2.75-inch CF flange was drilled with diameters of the holes large enough to fit several fibers. The fibers were inserted and epoxied in place using low out-gassing 3M Scotch-Weld 2216 epoxy.
@@
The DC and RF fields provided by the trap electrodes confine the newly ionized atom at the loading site when laser cooled. The RF field is driven at approximately 40 MHz to confine the ion radially with trapping frequencies near 3.1 MHz. DC fields confine ions axially with a trapping frequency of 1.02 MHz. RF power was monitored by a Texas Instruments LMH2120 power detector chip with a 1% capacitive pick-off circuit between the helical resonator and the trap RF input and is used in a servo to maintain constant RF power. The trapping pseudo-potential defined by the RF field holds ions approximately 50 m above the electrode surface, where all waveguide beams are expected to converge at their focuses at the interrogation sites. We measure an average motional quanta of quanta based on the fitting of Rabi oscillations [semenin2022determination] with a detection fidelity of 87%, which is comparable to previous research [Kwon:2023]. This low detection fidelity is due to the low numerical aperture and relatively high background scattering rate inherent in side-imaging.
@@
We note that the heating rate of the system is high compared to our previous work [Kwon:2023]. We think this is due to a combination of factors. For one, the ion could not be placed simultaneously on the RF-null at an optimal position for Doppler cooling while maintaining good overlap with the clock laser waveguide output. As a result of the ion being off the RF null, it is more susceptible to amplifier noise in the trap RF. We believe that this is due to imperfect trap fabrication.
@@
Additional heating could also be due to the charging of the waveguides by the 369 beam. As stated in previous sections, we noticed a correlation between the input 369 nm laser power and compensation values of the DC electrodes used to position the ions in the 369 and 435 beams. The charging effect also suggests that the ions have additional exposure to the dielectric surfaces and an implied increase in Johnson noise heating. This high heating rate prevents us from obtaining higher clock resolution with the Ramsey method and remains an improvement for future work.
@@
The trapped ions are moved to their target locations within a site by adjusting the DC potentials. Each location needs a specific field profile due to small fabrication errors and charging effects. Most importantly, these profiles are tuned to achieve identical Rabi dephasing and /2 pulse times. Therefore, we first used a dynamic solution to obtain precise control [Kwon:2023] and to optimize the compensation map for the trap. The dynamic solver provides a real-time calculation of the potential, which allows for easy compensation adjustment for various conditions. The dynamic solver allows for adjustment of a single trapping potential while maintaining the compensation fields at other trapping sites. For the actual clock algorithm, the DC voltages are compiled into a discrete set to reduce computational load and to increase and regularize the shuttling speed. This is critical for maintaining multi-ion occupancy for several hours. Consistent, fast shuttling reduces the accumulated motional quanta, which is important for high heating rate traps. The loss of an ion during shuttling becomes even more important as the number of ions increases. The ability to replenish lost ions depends on the mean loading rate being greater than the ion loss rate. The ion loss rate should scale as for ions and a single ion loss rate of . We addressed this by categorizing each ion loss event and mapping out a shuttling process that keeps the remaining ion sites stable. Additionally, up to half of the ensemble ions are shuttled simultaneously to minimize the number of required shuttling steps while not blocking the probe of clock ions. For the specific case described in this paper, four different shuttling scenarios cover the replacement of ions at the clock sites. All scenarios begin and end with all potential wells present at all sites. Fig.5 shows an example of some steps in the complied solution used in a single scenario. Step 0 is the state of all wells present at all sites. Each shuttling from one site to the other site takes less than a second. The number of shuttling scenarios is proportional to the number of ions, making this a conceivable path forward to managing large numbers of ions. The remaining limitations are the ion lifetimes, the tuning of individual sites, and the evolution of the stray field. These could be incorporated as part of a tuning procedure or possibly even dynamically adjusted for in more sophisticated algorithms. However, our stray field evolution did not occur on any meaningful timescale as long as 369 nm beam power remained nominal.
@@
We applied a similar side-imaging technique as in our previous research [Kwon:2023], but in a modified way. A 2-inch objective lens with a numerical aperture (NA) of 0.12 was placed at the side window to exclusively capture maximum ion fluorescence while minimizing light noise from waveguide emissions. We placed extra pinholes of 100 m diameter in front of the 1D array PMT (Hamamatsu H11659-200). Each pinhole was separated by 1 mm, equal to the pitch of the 1D PMT array. The image magnification was coarsely tuned by obtaining an image from a camera (Qimaging ROLERA em-c2). By coupling light to the integrated waveguides via fiber edge coupling, the stray light at the detectors was significantly decreased. With optimal alignment, we achieved a signal of kcps with kcps of noise, an SNR greater than 25. This is about five times better than the SNR from our previous work, allowing us to decrease the detection time to 3 ms.
Clock interrogation

### v3

@@ I Introduction
Optical atomic clocks, as time and frequency standards, offer unparalleled precision for applications in fundamental physics [16, 11, 28, 4, 17, 8], navigation [5], and telecommunications [14, 31]. Neutral atom optical lattice clocks are widely used platforms that provide high precision and accuracy [34, 3, 2, 18] but contain complex atom-atom interactions and require large high-power laser systems and atomic ovens. On the other hand, trapped-ion clocks [7, 9, 4], provide excellent isolation of individual ions with minimal disturbance of internal states. However, they suffer from high quantum projection noise as a result of the comparatively small number of ions. Although both platforms serve as excellent clocks, they both rely on bulky optical components, which limit their operation to laboratory environments.
The path to deploying these instruments in real-world applications lies in miniaturization and ruggedization [22, 5], a challenge for which integrated photonics provides a compelling solution. By replicating complex optical functionalities on scalable chip-based platforms [6, 32], this approach not only works to reduce the system’s footprint but also enables advanced multi-ion clock architectures that are less practical with conventional setups [27, 1]. Additionally, the integration of photonic components directly into ion trap chips is a key strategy for scaling these quantum systems. Replacing complex free-space optics with on-chip waveguides and splitters [33, 38, 13], grating couplers [20, 23, 19], and photonic control [10] provides robust and scalable delivery of the multiple laser wavelengths required for narrow transition control. These technologies are also applicable to trapped-ion quantum computing, notably the Quantum Charge-Coupled Device (QCCD) architecture [24, 25], which requires multiplexed operations and coherent control across many distinct trapping zones [21]. By providing a pathway to individually address dense ion arrays, integrated photonics are critical for moving beyond small-scale experiments, towards large-scale quantum platforms.
Recently demonstrated building blocks for this architecture include simultaneous multi-site optical addressing [13] and coherent transport of ions between zones [21], establishing the viability of complex on-chip control primitives. Although these advances are foundational, they have so far focused on individual capabilities rather than their combination into a unified, application-focused system. Complementing these efforts in ion trap integration, there is parallel progress in developing high-stability, chip-scale lasers for integrated, high-performance light sources [15] and integrated detectors for state readout [30, 26].
@@ II.1 Experimental setup
In previous demonstrations of integrated photonic ion traps, we utilized a free-space grating input coupling scheme [12, 13]. This presented challenges for both coupling efficiency and stability. Our current room-temperature system makes progress in both challenges by directly coupling an array of fibers to the trap. Single-mode, polarization-maintaining fibers in Pyrex v-groove arrays (VGAs) [37] were aligned with a matching array of tapered waveguides at the edge of the trap chip. The input for each beam is split on-trap via multi-mode interferometer (MMI) splitters into five waveguides, which then go to separate clock sites. Each site is addressable with four waveguide output grating windows that deliver light at 369 nm, 435 nm, 935 nm, and 760 nm. Fig.1(a) shows an illustration of the site layout on the chip. The optical loss in each waveguide path is primarily due to imperfect input coupling, material absorption, scattering during propagation, and diffraction inefficiency. The total loss for the clock beam (435 nm) exhibits a total loss of -27 dB. Although it is difficult to identify the contribution of each loss mechanism, measurements from the loop-back optimization suggest that most of the loss is due to the fiber-to-trap coupling.
For this work, we used the 435 nm E2 (clock) transition of . This additionally requires 369 nm for Doppler cooling and detection and 935 nm for repump and state preparation. A 760 nm beam can be used to pump the state to to rejoin the Doppler cooling cycle [35], however, in our systems it does not significantly improve the overall lifetime of the ions [13].
@@ II.2 Clock protocol for multi-ion ensemble operation
Fig.2(a) shows the clock pulse sequence used for interrogation. Clock ions are cooled and optically pumped. Then, the ensemble of ions is probed on the E2 transition at half its spectroscopic width on either side of the inferred center for two separate second-order integrators. Probing at the half-width ensures that fluorescence counts are on average present after sufficiently many interrogations if an ion is present. Additionally, this choice gives a large derivative in the count imbalance with respect to frequency error, the basis of the error signal. The count imbalance feeds into its respective integrator to maintain reference to the ion. Each integrator is independent of the other, probing the same transition on the same set of ions. Integrators follow closely as the ions are lost and reloaded at clock sites. For a single integrator, its feedback onto the local oscillator frequency is as follows. The ion is probed on the right and left sides of the transition one half-width from the inferred resonance. If the ion was shelved in the clock state and therefore below the threshold for detection, the value of the interrogation of that side () is ( is either or ). If the ion was not shelved in the clock state, visible to the detection cycle, and therefore above the threshold for detection, the value of the interrogation of that side is . The second-order integrator () is increased by the value of the first-order integrator (). The imbalance () is then added to the first-order of the integrator. Together, these steps are as follows:
@@
In addition to managing the integrator feedback, the clock algorithm uses the same measurement to track which ions are still present and whether an ion needs to be reloaded. To determine whether the ion is present, the sum of side interrogations is tracked. Both ions that are outside of the cooling cycle and physically lost from the trap are perceived as lost when the rolling average falls below a threshold. When an ion in a clock-zone is lost, this is reported to the shuttling software. This report occurs every 20 clock cycles. Each clock cycle involves four interrogations: one on each side of the resonance for each integrator. One clock cycle is the distance between the dashed green lines in Fig.2(a). The reporting rate is limited by the rate at which the receiving software can communicate consistently. The remaining ions are shuttled into vacancies, automatically filling the farthest sites from the loading site first. The new ion at the loading site is stored as a buffer to quickly replenish empty clock sites. An illustration of this can be seen in Fig.2(b). The lifetime of ions at an individual clock site was correlated with the distance from the loading site. This motivated filling the farthest sites first. One hypothesis for the cause of this effect is that the localized pressure associated with neutral flux and heated surfaces increases collisions, reducing neighboring clock site lifetimes. The ion lifetimes were approximately one minute during clock operation and 15 minutes when individual ions were only being trapped and cooled. Feedback onto the local oscillator occurs as long as any ions remain. Despite this short lifetime, we were able to keep the clock sites fully occupied with four ions more than 45% of the time. Furthermore, we operated the clock system continuously for more than two hours on multiple occasions, predominantly limited by algorithmic interlocks to abort the clock in cases where the algorithm was unable to determine if a laser had become unlocked, such as no ions in any sites for an extended period of time. This attests to the effectiveness of the automated ion reloading, shuttling, and clock-cycle reintegration capabilities of the system.
@@ II.3 Frequency stability of integrated photonics multi-ion clock
Fig.3 shows the performance of the integrated photonics 4-ion ensemble clock. The expected quantum projection noise (QPN) of a single integrator is calculated as follows [39]:
@@ II.4 Characterization of multi-site uniformity for future scalability
The direction and focal size of each waveguide output beam is lithographically defined. Before running the clock algorithm, the ions were verified to have similar counts, Rabi frequencies, and coherence times. The electrodes must be finely tuned to match these, as small local differences are hard to predict. We also observed a dependence of the compensation voltages on the 369 nm beam power. Because the timescale of the effect was hours, the small periods of the 369 nm beam being off for the clock transition excitation were inconsequential. Once established, the compensations would hold for several weeks. While we expect to eliminate this dependence in future traps by using conductive coatings on the output grating windows, the power drift was not an issue unless the laser power was changed by roughly a factor of 2 or more. Fig.4(a) shows a representative set of Rabi flops from an ion at each site. Measurements are modeled similarly to [29]. The ions are trapped in separate DC potentials 180 m apart, so the chain configuration terms can be dropped, resulting in the following equation for the probability of each ion being shelved in the clock state:
@@ III Conclusion
This trap suffers from high heating rates, as discussed in more detail in the Methods. Although this was not likely the limiting factor in clock performance, future systems with improved detection efficiency and ion lifetime would benefit from cooler atoms and the associated improved contrast and coherence time. One culprit for high heating rates could be noise due to the buildup of charge on the dielectric surfaces of the grating windows [36]. This could be alleviated by the inclusion of a conductive but transparent coating over the grating windows, such as indium-tin-oxide. In addition, side-band cooling was not implemented. Side-band cooling was time prohibitive, in part because of the low optical coupling from fiber-to-trap. Better optical coupling would allow for side-band cooling and increase both contrast and coherence time.
@@ Acknowledgments.
Data availability
Appendix A: Experimental system
@@
The experiment was conducted using a trap designed and fabricated at Sandia National Laboratories. Although this trap shares features with previous designs [12, 13], it uniquely incorporates a 15 multi-mode interferometer (MMI) waveguide splitter for simultaneous multi-ion clock interrogation. is a good species for deployable and robust systems due to its suppressed first-order Zeeman shift. The large wavelength range (369 nm to 935 nm) adds to the design complexity, as shorter wavelengths require unique waveguide designs and are more sensitive to fabrication tolerances and feature roughness. All waveguides, including 369nm and 435nm, were made with silicon nitride. This is made possible using a plasma-enhanced chemical vapor deposition (PECVD) nitride film which Sandia National Laboratories has optimized for low propagation loss at blue and UV wavelengths.
Fiber coupling was accomplished using a UHV vacuum-compatible glue attachments by Optelligent of v-groove arrays (VGA). Single-mode polarization-maintaining fibers of the appropriate core diameters (to ensure single-mode propagation at the appropriate wavelength) were placed in Pyrex v-groove arrays prior to attaching [31]. These were then aligned to a matching array of tapered waveguides on the edge of the trap chip. These tapered waveguide array edge-couplers are located transverse to the trap axis to minimize the scattered light from the coupling region into the readout optics. The attachment was optimized by maximizing the light that circulates through an additional on-chip loop-back waveguide. For a fiber feedthrough, a 2.75-inch CF flange was drilled with diameters of the holes large enough to fit several fibers. The fibers were inserted and epoxied in place using low out-gassing 3M Scotch-Weld 2216 epoxy.
@@
The DC and RF fields provided by the trap electrodes confine the newly ionized atom at the loading site when laser cooled. The RF field is driven at approximately 40 MHz to confine the ion radially with trapping frequencies near 3.1 MHz. DC fields confine ions axially with a trapping frequency of 1.02 MHz. RF power was monitored by a Texas Instruments LMH2120 power detector chip with a 1% capacitive pick-off circuit between the helical resonator and the trap RF input and is used in a servo to maintain constant RF power. The trapping pseudo-potential defined by the RF field holds ions approximately 50 m above the electrode surface, where all waveguide beams are expected to converge at their focuses at the interrogation sites. We measure an average motional quanta of quanta based on the fitting of Rabi oscillations [29] with a detection fidelity of 87%, which is comparable to previous research [13]. This low detection fidelity is due to the low numerical aperture and relatively high background scattering rate inherent in side-imaging.
@@
We note that the heating rate of the system is high compared to our previous work [13]. We think this is due to a combination of factors. For one, the ion could not be placed simultaneously on the RF-null at an optimal position for Doppler cooling while maintaining good overlap with the clock laser waveguide output. As a result of the ion being off the RF null, it is more susceptible to amplifier noise in the trap RF. We believe that this is due to imperfect trap fabrication.
@@
Additional heating could also be due to the charging of the waveguides by the 369 nm beam. As stated in previous sections, we noticed a correlation between the input 369 nm laser power and compensation values of the DC electrodes used to position the ions in the 369 nm and 435 nm beams. The charging effect also suggests that the ions have additional exposure to the dielectric surfaces and an implied increase in Johnson noise heating. This high heating rate prevents us from obtaining higher clock resolution with the Ramsey method and remains an improvement for future work.
@@
The trapped ions are moved to their target locations within a site by adjusting the DC potentials. Each location needs a specific field profile due to small fabrication errors and charging effects. Most importantly, these profiles are tuned to achieve identical Rabi dephasing and /2 pulse times. Therefore, we first used a dynamic solution to obtain precise control [13] and to optimize the compensation map for the trap. The dynamic solver provides a real-time calculation of the potential, which allows for easy compensation adjustment for various conditions. The dynamic solver allows for adjustment of a single trapping potential while maintaining the compensation fields at other trapping sites. For the actual clock algorithm, the DC voltages are compiled into a discrete set to reduce computational load and to increase and regularize the shuttling speed. This is critical for maintaining multi-ion occupancy for several hours. Consistent, fast shuttling reduces the accumulated motional quanta, which is important for high heating rate traps. The loss of an ion during shuttling becomes even more important as the number of ions increases. The ability to replenish lost ions depends on the mean loading rate being greater than the ion loss rate. The ion loss rate should scale as for ions and a single ion loss rate of . We addressed this by categorizing each ion loss event and mapping out a shuttling process that keeps the remaining ion sites stable. Additionally, up to half of the ensemble ions are shuttled simultaneously to minimize the number of required shuttling steps while not blocking the probe of clock ions. For the specific case described in this paper, four different shuttling scenarios cover the replacement of ions at the clock sites. All scenarios begin and end with all potential wells present at all sites. Fig.5 shows an example of some steps in the compiled solution used in a single scenario. Step 0 is the state of all wells present at all sites. Each shuttling from one site to the other site takes less than a second. The number of shuttling scenarios is proportional to the number of ions, making this a conceivable path forward to managing large numbers of ions. The remaining limitations are the ion lifetimes, the tuning of individual sites, and the evolution of the stray field. These could be incorporated as part of a tuning procedure or possibly even dynamically adjusted for in more sophisticated algorithms. However, our stray field evolution did not occur on any meaningful timescale as long as 369 nm beam power remained nominal.
@@
We applied a similar side-imaging technique as in our previous research [13], but in a modified way. A 2-inch objective lens with a numerical aperture (NA) of 0.12 was placed at the side window to exclusively capture maximum ion fluorescence while minimizing light noise from waveguide emissions. We placed extra pinholes of 100 m diameter in front of the 1D array PMT (Hamamatsu H11659-200). Each pinhole was separated by 1 mm, equal to the pitch of the 1D PMT array. The image magnification was coarsely tuned by obtaining an image from a camera (Qimaging ROLERA em-c2). By coupling light to the integrated waveguides via fiber edge coupling, the stray light at the detectors was significantly decreased. With optimal alignment, we achieved a signal of kcps with kcps of noise, an SNR greater than 25. This is about five times better than the SNR from our previous work, allowing us to decrease the detection time to 3 ms.
Appendix B: Clock interrogation